While exploring a cave, a spelunker starts at the entrance and moves the following distances: 75.0m north, 300m east, 210m at angle of 30.0 degrees north of east, and 150m south. Find the resultant displacement from the cave entrance.
1
Expert's answer
2012-09-28T08:44:50-0400
Assuming north and east directions as positive, getting:
N = 75.0 + 210 *
sin30 - 150 = 30 m
E = 300.0 + 210 * cos30 = 481.9 m
So, the resultant
displacement L = sqrt( N^2 + E^2 ) = sqrt(900 +232194) = 482.8 m